Ricky and Krystal have a new grandson. How much money should they invest now so that he will have $63,000 for his college education in 18 years? The money is invested at 5.6% compounded quarterly.

Answer by Boreal(15235)

P=Po(1+r/n)^nt
63000=Po(1+.056/4)^72, since there will be 72 quarters in 18 years
63000=Po(1.014^72)
Po=63000/2.72 round at end
=$23153.03